# Implicit Reasoning in Deep Time Series Forecasting

### v3

@@ 2 Related work
Prior research has explored implicit reasoning in language models [2, 1, 30, 25, 32]. Implicit reasoning is often assessed through tasks that require models to apply knowledge learned during training to new test instances. One common form is composition, where models must chain multiple facts to answer a question [25, 32, 30]. Other forms of implicit reasoning explored include comparison and inverse search [2, 1]. Comparison involves models making judgments about unseen or altered data, such as comparing attributes of entities observed during training. Inverse search tests the ability to retrieve information in the reverse order of the training task. For instance, identifying an entity based on its attributes when the model was trained for the reverse task. More information on related work is provided in Appendix A.1. No prior research has conducted controlled experiments in time series forecasting to evaluate implicit reasoning on OOD data. Our study addresses this gap by introducing a novel framework that aligns LLM research with time series models, offering insights into optimal architectures for future TSFM development.
@@ 3.1 Implicit Reasoning Tasks
We assess the implicit reasoning capabilities of models, or their ability to apply and manipulate internalized patterns in novel situations, through experiments inspired by language model research [2, 1, 25, 30, 32]. We carefully design experiments to test 3 forms of implicit reasoning: composition, comparison, and inverse search. We follow the notation from [25, 32], where ‘facts’ are formatted as (subject, relation, object). For continuous data, we adapt this to (input, function, output), with defining the relationship between the function , input , and output . We consider two sets of basis functions: and where is the set of all constant signals with linear trends and is the set of all sinusoidal signals with amplitude , frequency , and baseline shift . We also define a set of functions consisting of compositions of basis functions: . Given a set of data-generating functions and their parameters observed during training, which are considered in-distribution (ID), denoted as , we assess a model’s ability to generalize to other data-generating functions observed only at inference, which we consider OOD, denoted as .
@@
For models trained on signals generated from basis functions considered ID, we assess whether they can accurately forecast compositions of these signals that are considered OOD. Specifically, we define the datasets as follows: . Consider a specific example:
@@
Comparison latent reasoning refers to a model’s ability to make comparative judgments on unseen or altered data. For models trained on data generated from functions with parameters considered ID, we evaluate whether these models can generalize to data generated by the same functions with larger or smaller parameter values that are considered OOD. Specifically, we define the datasets as follows: . Consider a specific example:
@@
We evaluate whether models can ‘disentangle’ information through the inverse search task, which requires that models trained on composite signals are able to infer decompositions of functions to accurately forecast each of the OOD target signals . Specifically, we define the datasets as follows: . Consider a specific example:
@@ 3.2 Data
| Task | Dist. | Function | Parameters |
@@ A.1 Extended Related Work
Prior research has explored implicit reasoning in language models [2, 1, 30, 25, 32]. Implicit reasoning is often assessed through tasks that require models to apply knowledge learned during training to new test instances. One common form is composition, or multi-hop reasoning, where models must chain multiple facts to answer a question [25, 32, 30]. Other forms of implicit reasoning explored include comparison and inverse search [2, 1]. Comparison involves models making judgments about unseen or altered data, such as comparing attributes of entities from training. Inverse search tests the ability to retrieve information in the reverse order of the training task. An example of an inverse order task is identifying an entity based on its attributes when the model was trained for the reverse task. Allen-Zhu et al. [1] found that generative models struggle with inverse search unless pretrained for it specifically. [1].

### v4

@@ 2 Related work
Prior research has explored implicit reasoning in language models [1, 2, 25, 30, 32]. Implicit reasoning is often assessed through tasks that require models to apply knowledge learned during training to new test instances. One common form is composition, where models must chain multiple facts to answer a question [25, 30, 32]. Other forms of implicit reasoning explored include comparison and inverse search [1, 25]. Comparison involves models evaluating two or more entities to make judgments, such as determining whether the attribute value of one entity is greater or smaller than that of another. Inverse search tests a model’s ability to generate predictions in the reverse order of the training task. For example, this could involve applying the model to identify an entity based on its attributes when it was originally trained to predict the attributes of entities. More information on related work is provided in Appendix A.1. No prior research has conducted controlled experiments in time series forecasting to evaluate implicit reasoning on OOD data. Our study addresses this gap by introducing a novel framework that aligns LLM research with time series models, offering insights into optimal architectures for future TSFM development.
@@ 3.1 Implicit Reasoning Tasks
We assess the implicit reasoning capabilities of models, or their ability to apply and manipulate internalized patterns in novel situations, through experiments inspired by language model research [1, 2, 25, 30, 32]. We carefully design experiments to test 3 forms of implicit reasoning: composition, comparison, and inverse search. We follow the notation from [25, 32], where ‘facts’ are formatted as (subject, relation, object). For continuous data, we adapt this to (input, function, output), with defining the relationship between the function , input , and output . We consider two sets of basis functions: and where is the set of all constant signals with linear trends and is the set of all sinusoidal signals with amplitude , frequency , and baseline shift . We also define a set of functions consisting of compositions of basis functions: . For a set of data-generating functions and their parameters included in the training set and referred to as in-distribution (ID) or , we assess a model’s ability to generalize to other disjoint sets of data-generating functions observed only at inference, denoted as , which are considered OOD.
@@
For models trained on signals generated from basis functions considered ID, we assess whether they can accurately forecast signal compositions that are considered OOD. Specifically, we define the datasets as follows: . Consider a specific example:
@@
Comparison refers to a model’s ability to make comparative judgments about the attributes or parameter values of facts [25]. For models trained on data generated from functions considered ID, we evaluate whether these models can generalize to the same data-generating function with larger or smaller parameter values that are considered OOD. Specifically, we define the datasets as follows: . Consider a specific example:
@@
We evaluate whether models can decompose functions through the inverse search task. For models trained on signal compositions, we assess their ability to accurately forecast individual basis functions comprising the signal, which are considered OOD. We define the datasets as follows: . Consider a specific example:
@@ 3.2 Data
| Task | Dataset | Function | Parameters |
@@ A.1 Extended Related Work
Prior research has explored implicit reasoning in language models [1, 2, 25, 30, 32]. Implicit reasoning is often assessed through tasks that require models to apply knowledge learned during training to new test instances. One common form is composition, where models must chain multiple facts to answer a question [25, 30, 32]. Other forms of implicit reasoning explored include comparison and inverse search [1, 25]. Comparison involves models evaluating two or more entities to make judgments, such as determining whether the attribute value of one entity is greater or smaller than that of another. Inverse search tests a model’s ability to generate predictions in the reverse order of the training task. For example, this could involve applying the model to identify an entity based on its attributes when it was originally trained to predict the attributes of entities. Allen-Zhu et al. found that generative models struggle with inverse search unless pretrained for it specifically [1].